slope of 2 and y-intercept of -4 turned into an equation
Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

slope of 2 and y-intercept of -4 turned into an equation
--------
It's y = mx + b where m = slope and b is the y-intercept

Answer by ikleyn(52847)

.
    y - (-4) = 2x,


or, equivalently,


    y + 4 = 2x,


or


    2x - y = 4.


Choose any form of an equation.